OpenStudy (anonymous):

I am not sure how to work out these equations... sqrt(4x-7)=3-x

OpenStudy (anonymous):

square both sides as step one, to get rid of the radical on the left

OpenStudy (anonymous):

\[\sqrt{4x-7}=3-x\] \[\sqrt{4x-7}^2=(3-x)^2\] \[4x-7=9-6x+x^2\] (dont forget how to square correctly \((3-x)^2=(3-x)(3-x)=9-3x-3x+x^2=9-6x+x^2\)

OpenStudy (anonymous):

your next job is to solve \[4x-7=9-6x+x^2\] do you know how to do that?